# PRIMA.CPP: Speeding Up 70B-Scale LLM Inference on Low-Resource Everyday Home Clusters

### v1

@@ Abstract
Emergency of DeepSeek R1 and QwQ 32B have broken through performance barriers for running frontier large language models (LLMs) on home devices. While consumer hardware is getting stronger and model quantization is improving, existing end-side solutions still demand GPU clusters, large RAM/VRAM, and high bandwidth, far beyond what a common home cluster can handle. This paper introduces prima.cpp, a distributed inference system that runs 70B-scale models on everyday home devices using a mix of CPU/GPU, low RAM/VRAM, Wi-Fi, and cross-platform support. It uses mmap to manage model weights and introduces piped-ring parallelism with prefetching to hide disk loading. By modeling heterogeneity in computation, communication, disk, memory (and its management behavior), and OS, it optimally assigns model layers to each device’s CPU and GPU, further reducing token latency. An elegant algorithm named Halda is proposed to solve this NP-hard assignment problem. We evaluate prima.cpp on a common four-node home cluster. It outperforms llama.cpp, exo, and dllama on 30B+ models while keeping memory pressure below 6%. This brings frontier 30B-70B models, such as Llama 3, DeepSeek R1, Qwen 2.5, and QwQ to home assistants, making advanced AI truly accessible to individuals. The code is open source and available at https://github.com/Lizonghang/prima.cpp.
@@ 1 Introduction
Current large language models (LLMs) are mostly cloud-based, but DeepSeek (DeepSeek-AI, 2025) changes this. Its R1 series, spanning from 1.5B to 70B, bring small models closer to frontier performance. Its 70B version even surpasses cloud-based models like GPT-4o and Claude 3.5 Sonnet. This has sparked interest in deploying LLMs locally on users’ own devices, especially with the explosion of open-source LLMs. However, limited by weak chips and small RAM, user devices struggle to run anything beyond 10B, even with 4-bit quantization. For example, running Qwen 2.5-14B (Q4K) on a Mac M1 with 8 GiB RAM takes a staggering 10 seconds per token. As a result, most end-side LLM systems prefer smaller models, e.g., run 7B models on phones and browsers (MLC, 2023; Lugaresi et al., 2019) and 3.8B models on an Android device (Ghorbani, 2024).
That’s not too bad, users have many devices, e.g., laptops, desktops, phones, and tablets. Some PCs have high-end GPUs like the NVIDIA 20/30/40 series, and Mac M-series have Apple Metal GPUs. By pooling the computing power and memory of home devices, it’s possible to run larger models. In pipeline parallelism, the model is split into segments and assigned to devices. Exo (Exo, 2024) distributes model layers based on memory size, devices with more RAM/VRAM handle more layers. While this prevents OOM, it can slow speed if high-memory devices have weak CPUs/GPUs. This raises a question: How should model layers be assigned? A simple alternative is to assign model layers based on computing power, with stronger devices handling more layers, but also with high OOM risk. Ye et al. (2024) offers an idea: first, assign model layers based on computing power, then migrate layers from OOM devices to those with free memory. However, this also slows down inference, as weaker CPUs handle more layers.
So new questions arise: Can stronger devices handle these overloaded layers instead? If we require the cluster memory to meet the model’s needs (Ye et al., 2024; Exo, 2024; Tadych, 2024; Zhang et al., 2025; Lee et al., 2024; Zhao et al., 2023; Zhang et al., 2024), the answer is no, as stronger devices have reached their OOM limits. But what if we relax this requirement? For example, we can free up processed layers on stronger devices and load the extra layers to continue computation. With a fast SSD, this could be more efficient than using weak-CPU devices. This adds complexity: whose disks are faster than weak CPUs? And how many extra layers can be migrated to maximize inference speed? This is tricky because disk I/O introduces a new bottleneck: disk loading latency, which depends on data size to be (re-)load and disk read speed. These two factors vary with hardware and OS differences. For example, macOS with Metal reclaims memory more aggressively than Linux, causing more reloads, while Linux optimizes sequential reads, making reloading faster than on macOS. This heterogeneity makes disk loading latency hard to estimate.
This paper introduces a device profiler to capture system heterogeneity across computation, memory, disk, communication, and OS. We mathematically model token latency from these factors and propose Halda, an efficient algorithm to solve this layer-to-device assignment (LDA) problem. Halda provides an optimal workload distribution across CPU and GPU (if available) for each device. It determines how many model layers each device should handle and how to allocate them between CPU and GPU. To hide disk loading, we design piped-ring parallelism, where devices form a ring and pass their results to the next for subsequent processing. Unlike existing systems, in our design, devices can complete a token with multiple rounds, and with prefetching, we overlap disk loading with other devices to hide latency. We call it prima.cpp, as it uses piped-ring parallelism and builts on llama.cpp (Gerganov, 2024). The main contributions are summarized as follows:
  • We propose prima.cpp, a distributed inference system designed for low-resource home clusters. It uses mmap to lazily load model weights, and piped-ring parallelism with prefetching to hide disk latency. It prevents OOM and reduces token latency.
  • We model the LDA problem and develop a device profiler to capture the system heterogeneity. It minimizes token latency by modeling delays in computation, memory access, disk loading, and communication, while optimizing the use of RAM and VRAM.
  • We propose the Halda algorithm to solve the LDA problem. Halda breaks the NP-hard problem into a set of simple ILPs, so we can find the optimal solution in polynomial time.
  • We implement prima.cpp with 20K LoC modifications. Evaluation on a real home cluster shows that prima.cpp is 15× faster than llama.cpp on 70B models, with memory pressure below 6% per device. It also surpasses distributed alternatives like exo (Exo, 2024) and dllama (Tadych, 2024) in both speed and memory efficiency across all 7B-72B models.
In our experiments, a small, heterogeneous, and budget-friendly home cluster (2 laptops, 1 desktop, 1 phone) was used. Our prima.cpp achieves $\sim$600 milliseconds per token and a time-to-first-token (TTFT) below 2 seconds for a 70B model, making it accessible for voice chat apps like home Siri. For 45B-70B models, the speed matches those of audiobook apps. It now supports hot models including Llama 3, Qwen 2.5, QwQ, and DeepSeek R1 (distilled versions).
Table 1: Comparison of distributed LLM inference systems at the edge.
Type1 Backends2 Mem3 Quantization Mem Pressure4 Speed Heterogeneity5
dllama (Tadych, 2024) TP CPU RAM Q4 Critical Slow
Zhang et al. (2025) TP CPU RAM FP32 Critical Slow
Hepti (Lee et al., 2024) TP CPU RAM FP32 Critical Slow
Galaxy (Ye et al., 2024) TP+SP CPU / GPU RAM / VRAM FP32 Critical Slow
TPI-LLM (Li et al., 2024) TP CPU RAM FP32 Medium Slow
exo (Exo, 2024) PP CPU / GPU RAM / VRAM Q4+FP32 Critical Slow
LinguaLinked (Zhao et al., 2023) PP CPU RAM Q8 / FP32 Critical Slow
EdgeShard (Zhang et al., 2024) PP CPU / GPU RAM / VRAM FP32 Critical Fast
prima.cpp (ours) PRP CPU & GPU RAM & VRAM Q4 / IQ1 Low Fast
  • 1
    “TP" is tensor parallelism, “PP" is pipeline parallelism, “SP" is sequential parallelism, “PRP" is piped-ring parallelism.
  • 2
    Backends simultaneously used on one device. “CPU / GPU" indicates only one backend is used at a time on a device, and if a GPU is available, only the GPU will be used. “CPU & GPU" indicates that both CPU and GPU are used if a GPU is available.
  • 3
    Memory simultaneously used on one device. The definitions of “RAM / VRAM" and “RAM & VRAM" are similar to the above.
  • 4
    Memory pressure when the cluster is just enough to run the model. Low: memory can be reclaimed normally. Medium: Potential for page swapping. Critical: High pressure, likely to freeze or be killed.
  • 5
    Whether be optimized for heterogeneous devices.
@@ 2 Related Work
On-device LLM systems. Most of these systems are limited to small models. MLC-LLM (MLC, 2023) and MediaPipe (Lugaresi et al., 2019) bring 7B models to mobile phones and browsers. PocketPal AI (Ghorbani, 2024), which runs on Android using llama.cpp (Gerganov, 2024), supports models up to 3.8B. Some efforts push for larger models, like AirLLM (Li, 2023), which loads only needed layers to save memory but at the cost of speed. Others turn to high-end hardware, such as the Apple M2 Ultra (192 GiB RAM) for a 65B model or kTransformers (kvcache ai, 2025) (382 GiB RAM) for a 671B model ($\sim$75 GiB for 70B). These setups (large RAM, advanced CPUs with specific instruction sets) go far beyond common home devices and are inaccessible to most users.
Distributed LLM systems at the user edge. Distributed systems break the limits of a single device. We focus on edge LLM inference and categorize recent efforts into tensor and pipeline parallelism.
@@
- Tensor Parallelism. Tensor parallelism splits weight tensors (e.g., attention heads and FFNs) across devices to share the load (Shoeybi et al., 2019). To reduce all-reduce costs, dllama (Tadych, 2024) uses USB4 and Thunderbolt 5 for fast connections, and (Zhang et al., 2025) uses wireless analog superposition to perform all-reduce over the air. Due to device heterogeneity, Hepti (Lee et al., 2024) optimizes workload partitioning with three slicing strategies for different memory budgets, and Galaxy (Ye et al., 2024) prioritizes compute power first, then memory, to maximize speed and avoid OOM. These systems load the full model into cluster memory. With limited total memory, only small models can run. TPI-LLM (Li et al., 2024) loads model layers on demand and hides disk loading with prefetching. This allows poor devices with only 4 GiB RAM to run a 70B model, but it’s still slow.
@@
- Pipeline Parallelism. Due to Wi-Fi’s high latency, pipeline parallelism is better for home clusters as its fewer P2P communication. Exo (2024); Zhao et al. (2023); Zhang et al. (2024) split the model into segments and assign them to devices based on memory, compute, and network conditions. Each device computes its segment and passes the result to the next, until the last device outputs the next token. Exo (2024) partitions model segments based on memory ratio; LinguaLinked (Zhao et al., 2023) uses linear optimization to solve the device assignment problem; and EdgeShard (Zhang et al., 2024) uses dynamic programming. Some of them are efficient, but rely on specialized devices like Jetson AGX/Nano, which are not common in household.
These distributed systems, except TPI-LLM, have severe limitation: they require cluster memory to meet the model’s needs. When cluster memory is limited, only small models can run. However, users have few devices with less mem_available (could be much less than mem_total), and allocating too much for a LLM app can freeze devices and degrade user experience. Additionally, as shown in Table 1, some systems lack GPU support, while others don’t use CPU offloading. This restricts devices to either RAM or VRAM and further limiting cluster memory. Most also skip quantization and reside model weights in memory, making OOM highly likely for larger models.
Prima.cpp addresses these issues with four key features: (a) support disk offloading to allow cluster memory to be less than the model’s needs. While this causes disk latency, prima.cpp minimizes it through optimal workload distribution and piped-ring parallelism with prefetching; (b) support GPU & CPU offloading to combine RAM and VRAM into cluster memory, with optimal workload assignment to alleviate CPU slowdowns; (c) support quantization like Q4K (used by default) and IQ1 to cut memory needs; and (d) use mmap to cache model weights, so the OS can free them as needed.
3 Prima.cpp: Parallel Architecture and Scheduler Design
We first introduce the piped-ring parallelism with prefetching adopted by prima.cpp in Section 3.1. Section 3.2 then mathematically models the layer-to-device assignment problem and Section 3.3 proposes a polynomial-time algorithm to solve it, so that the inference latency is minimized.
3.1 Piped-ring Parallelism with Prefetching
Pipeline parallelism is common in distributed inference, where the model is split into layer-based segments and assigned to a chain of devices. Each device computes its segment and passes the result to the next, this process continues until the last device outputs the next token. Exo (2024); Zhang et al. (2024); Zhao et al. (2023) take this a step further by linking the last device back to the first, forming a ring. This allows input and output to be processed on the head device, providing enhanced privacy.
Pipeline parallelism works well for batch inference when resources are abundant, but it’s not a good fit for resource-limited home clusters due to: (a) home users have few devices with low specs (see Table 2); (b) LLM apps should use mem_available, not mem_total, or other apps will slow down, degrading user experience; and (c) batch inference demands more RAM. These limitations make it hard to gather enough devices to meet the memory needs of larger models. In this study, we focus on single-request inference to tackle the immediate challenge of fast inferencing 70B models in resource-scarce clusters. Batch inference can be developed once this foundation is solid.
@@ 3.1 Piped-ring Parallelism with Prefetching
Figure 1: Piped-ring parallelism. In this case, 6 devices handle a 36-layer model. With a layer window size of 2, the model is splitted into 18 segments, which are assigned to 6 devices in a ring order, so each device needs 3 rounds to predict one token.
Given this setup, the drawback of pipeline becomes apparent: when a device is running, it loads required layers (due to limited mem_available, we swap layers from disk instead of keeping a full model) and computes them, meanwhile, others sit idle. To improve this, we use prefetching.
Pipeline parallelism with prefetching. Prefetching loads layers in advance, ensuring that when computing starts, the required layers are already (or partially) in place. This overlaps disk loading with ongoing operations on other devices, so part of the latency can be hidden. We use mmap advice to achieve this: once a device finishes computing, we mark upcoming layers as WILLNEED to advise the OS to prefetch them in the background when the load permits, so it does not block the main process. It continues even after computing starts until page faults occur.
This vanilla approach has a flaw: if the current device has a high disk speed, or other devices run for a long time, prefetching may load too many layers, exceeding its available memory. As a result, later-loaded layers will overwrite those prefetched earlier. We call this effect “prefetch-release" (see Appendix A.1), which results in all layers being loaded twice, adding unnecessary disk overhead without any benefit from prefetching. To address this, we propose the piped-ring parallelism.
Piped-ring parallelism with prefetching. In piped-ring parallelism, devices are connected in a ring structure, but unlike Zhang et al. (2024); Exo (2024); Zhao et al. (2023), we can run multiple rounds to predict one token. We define layer window size as the number of model layers a device should handle in each round. For example, in Figure 1, each device takes 3 rounds to predict one token. In each round, device 1 should handle 2 model layers, so its window size is 2. Layer window sizes vary by device capability—stronger devices have larger windows. However, finding the optimal window size setup is complex. In section 3.2, we mathematically model this problem, followed by an efficient algorithm in section 3.3 to solve it. By setting the layer window size small, we ensure the model layers stays within memory limits, avoiding “prefetch-release" during prefetching. We visualize this magic and explain how disk loading is overlapped in Appendix A.2.

### v2

# Prima.cpp: Fast 30-70B LLM Inference on Heterogeneous and Low-Resource Home Clusters

@@ Abstract
On-device inference offers privacy, offline use, and instant response, but consumer hardware restricts large language models (LLMs) to low throughput and capability. To overcome this challenge, we present prima.cpp, a distributed on-device inference system that runs 30-70B LLMs on consumer home clusters with mixed CPUs/GPUs, insufficient RAM/VRAM, slow disks, Wi-Fi links, and heterogeneous OSs. We introduce pipelined-ring parallelism (PRP) to overlap disk I/O with compute and communication, and address the prefetch-release conflict in mmap-based offloading. We further propose Halda, a heterogeneity-aware scheduler that co-optimizes per-device CPU/GPU workloads and device selection under RAM/VRAM constraints. On four consumer home devices, a 70B model reaches 674 ms/token TPOT with <6% memory pressure, and a 32B model with speculative decoding achieves 26 tokens/s. Compared with llama.cpp, exo, and dllama, our proposed prima.cpp achieves 5-17× lower TPOT, supports fine-grained model sizes from 8B to 70B, ensures broader cross-OS and quantization compatibility, and remains OOM-free, while also being Wi-Fi tolerant, privacy-preserving, and hardware-independent. The code is available at https://gitee.com/zonghang-li/prima.cpp.
@@ 1 Introduction
As a representative of next-generation AI, embodied AI (Gupta et al., 2021) is acutely sensitive to privacy protection, stable connectivity, and long-term financial costs. Home surveillance, always-listening assistants, and companion robots cannot send raw interactions to the cloud. Cloud services also suffer from frequent network failures, queuing, timeouts, and their costs scale linearly with token usage and service time. Consequently, embodied AI is moving from the cloud to the device.
However, on-device inference is constrained by modest chips and small RAM, struggling to run models beyond 8B (MLC, 2025; Lugaresi et al., 2019; Ghorbani, 2025), yet reliable long-term planning and tool use require 32B or more. To support larger models, disk offloading helps capacity but at the cost of speed (Gerganov, 2025; Li, 2023), e.g., Qwen 2.5-14B Q4K on an 8 GiB Mac M1 laptop takes 10 s/token with llama.cpp. Recent efforts use dedicated devices (e.g., Jetson, Mac Studio) for larger models and faster speed (Ye et al., 2024; 2025), but they are too expensive for most households. This motivates our goal: Can we deliver fast 30-70B inference on user-owned consumer devices that meet embodied AI’s demands summarized in Table 1?
Distributed inference may be the only solution to improve model size and speed while preserving the original outputs. Users usually own multiple devices, e.g., laptops, PCs, phones, and tablets. Some laptops and PCs have low-end GPUs like NVIDIA 20-/30-/40-series, and some Macs have Apple Silicon GPUs. By pooling the computing and memory of home devices, we can run larger models.
Despite recent progress, four limitations remain: (a) Existing systems require sufficient aggregated RAM/VRAM to hold the full model (Ye et al., 2024; Exo, 2025; Tadych, 2025; Zhang et al., 2025; Lee et al., 2024; Zhao et al., 2023; Zhang et al., 2024), driving up hardware costs and restricting model size. (b) Disk offloading slows inference, causing time-per-output-token (TPOT) of tens of seconds (Li et al., 2025). (c) Layer partitioning relies on the strong assumption in (a) and overlooks heterogeneity in OS-specific memory reclamation and disk access. (d) All devices are assumed to be necessary, even though removing slow ones could improve speed. These limitations raise two questions that motivate our design: (Q1) How can memory constraints be relaxed to run larger models? If disk offloading is required, how can disk latency be hidden? (Q2) Given Q1, how can we design heterogeneity-aware layer partitioning and identify bottleneck devices?
We propose prima.cpp, the first distributed inference system for consumer home clusters, with mixed CPUs/GPUs, insufficient RAM/VRAM, slow disks, Wi-Fi links, heterogeneous OSs, and capable of running 30-70B models at practical speed. To address Q1, Section 3.1 proposes pipelined-ring parallelism with prefetching to overlap disk latency and resolve the prefetch-release conflict in mmap. To address Q2, Section 3.2 models heterogeneity in compute, communication, memory, and OS-specific memory reclamation and disk optimizations. Section 3.3 develops Halda to find the optimal layer partitioning and select the best-performing devices. Section 4 presents experiments on real home clusters and shows that, even under limited memory, a 70B model runs locally at 674 ms/token with <6% memory pressure. With speculative decoding (Leviathan et al., 2023), a 32B model achieves 26 tokens/s, which is fast enough and paves the way for household embodied AI.
As summarized in Table 1, prima.cpp runs on existing, free devices; all data stays local; no queueing or timeouts; offline; and supports models up to 70B. To the best of our knowledge, prima.cpp is the first on-device system to deliver practical performance for 30-70B models in such constrained environments, without relying on specialized hardware (e.g., Jetson, NPUs) or altering model outputs.
Table 1: Comparison of cloud and local LLM deployments. TPS1 indicates the per-request token rate, TPS2 denotes TPS values measured across 32-70B models. OA denotes open-access.
Cost (\$) Privacy Speed (TPS) Queue Net Support Model
Cloud App < 100 10-701 Specific models
API < 100 10-701 Specific models
Dedicated < 1,000 10-352 Any OA model
Local Dedicated > 1,000 4-172 Any OA model ($\leq$70B)
Consumer Free 2-262 Any OA model ($\leq$70B)
@@ 2 Related Work
On-device LLM systems. Most of them run on a resource-constrained device and can only handle small models. MLC-LLM (MLC, 2025) and MediaPipe (Lugaresi et al., 2019) bring 7B models to mobile phones and browsers. PocketPal AI (Ghorbani, 2025), which runs on Android using llama.cpp (Gerganov, 2025), supports models up to 3.8B. Some efforts push for larger models, like AirLLM (Li, 2023), which loads only the needed layers to save memory but at the cost of speed. Others turn to high-end hardware, such as the Apple M2 Ultra (192 GiB RAM) for a 65B model or kTransformers (kvcache ai, 2025) (382 GiB RAM) for a 671B model ($\sim$75 GiB for 70B). Like kTransformers, HeteGen (Zhao et al., 2024) also collaborates with the CPU and GPU on one device to accelerate inference. These setups (large RAM, advanced CPUs with specific instruction sets, dedicated hardware) go far beyond common home devices and are inaccessible to most households.
Distributed on-device LLM systems. Such distributed systems follow two main parallelism paradigms: tensor parallelism and pipeline parallelism.
@@
- Tensor Parallelism (TP). TP splits tensors across devices to share the load (Shoeybi et al., 2019). To speed up all-reduce, dllama (Tadych, 2025) uses USB4 and Thunderbolt 5 for fast connections, and AirInfer (Zhang et al., 2025) uses wireless analog superposition to perform all-reduce over the air. Due to device heterogeneity, Hepti (Lee et al., 2024) optimizes workload partitioning with three slicing strategies for different memory budgets, and Galaxy (Ye et al., 2024) prioritizes compute power first, then memory, to maximize speed and avoid OOM. These systems reside the full model in the aggregate memory. With limited aggregate memory, only small models can run. TPI-LLM (Li et al., 2025) loads model layers on demand and hides disk loading with prefetching. This enables low-end devices with only 4 GiB of RAM to run a 70B model, but at 30 s/token, making it impractical.
@@
- Pipeline Parallelism (PP). Due to Wi-Fi’s high latency, pipeline parallelism becomes more suitable for home clusters as it requires less P2P communication. Exo (2025); Zhao et al. (2023); Zhang et al. (2024) split the model into segments and assign them to devices based on memory, compute, and network conditions. Each device computes its segment and passes the result to the next, until the last device outputs the next token. Exo (2025) partitions model segments based on memory ratio; LinguaLinked (Zhao et al., 2023) uses linear optimization to solve the device assignment problem; and EdgeShard (Zhang et al., 2024) uses dynamic programming.
However, these systems either require dedicated hardware (e.g., Jetson AGX/Nano) or sufficient aggregate memory to reside the full model, support only CPU or GPU backends, overlook heterogeneity (especially in disk offloading), and impose high memory pressure. Table 2 summarizes their features. These limitations raise hardware costs, restrict use to small models, cause slow inference and device freezes, and discourage users from deploying LLMs at home.
Instead, prima.cpp runs on consumer devices, supports disk offloading to run larger models, and uses pipelined-ring parallelism (PRP) to overlap disk latency. It runs on both GPUs and CPUs, combines RAM and VRAM, and models system heterogeneity to optimize GPU-CPU layer partitioning per device. Moreover, it stores model weights in the OS page cache, allowing the OS to reclaim memory as needed to preserve user experience. These features distinguish prima.cpp from existing systems.
Table 2: Comparison of distributed on-device LLM systems. (Abbr.: Quantization, Heterogeneity)
Type Backends Mem Quant. Mem Stress Speed Hete.
dllama TP CPU RAM Q4 Critical Slow
AirInfer TP CPU RAM FP32 Critical Slow
Hepti TP CPU RAM FP32 Critical Slow
Galaxy TP+SP CPU/GPU RAM/VRAM FP32 Critical Slow
TPI-LLM TP CPU RAM FP32 Medium Slow
exo $P$ CPU/GPU RAM/VRAM Q4+FP32 Critical Slow
LinguaLinked $P$ CPU RAM Q8/FP32 Critical Slow
EdgeShard $P$ CPU/GPU RAM/VRAM FP32 Critical Fast
prima.cpp PRP CPU&GPU RAM&VRAM Q4/IQ1 Low Fast
3 PRIMA.CPP: use Pipelined-RIng parallelism in llaMA.CPP
3.1 Pipelined-Ring Parallelism with Prefetching
PP is effective for batched inference when aggregated memory is abundant; however, this prerequisite rarely holds in households. In such cases, on-device LLMs typically serve very few users at low frequency, often a single request at a time [^1] We target a single request, but this can extend to batches via dynamic batching, which we have implemented., which prevents mini-batching and leaves significant pipeline bubbles. Besides, home devices are few and have limited available memory [^2] We cannot occupy all memory, as this may disrupt other apps (e.g., TikTok) and cause users to kill the LLM., making it difficult for users to build a cluster that can hold the full model in memory.
To relax memory constraints, we extend PP with mmap, building on llama.cpp (Gerganov, 2025): when computation requires the layers, mmap loads them from external storage on demand, and lets the OS evict them under memory pressure, so a small-memory cluster can run larger models. To hide disk loading latency, we employ prefetching, allowing devices to preload the next model segment.
This naive design supports larger models with low memory pressure but suffers from slow speed, because prefetching will fail due to prefetch-release conflict: if disk reads are fast, later-loaded layers will evict earlier prefetched layers from page cache, so when computation begins, the needed layers are no longer cached (see Appendix A.1 for details). This triggers page faults and reloads, negating the benefit of prefetching, and pipeline bubbles remain significant (see Fig. 6 d,e in Appendix A.2).
Pipelined-ring parallelism (PRP) with prefetching. To address the prefetch-release conflict, we further propose PRP, which connects devices end-to-end in a ring and runs multiple rounds to predict one token. In each round, devices prefetch different layer segments from the disk, which overlap with other devices’ ongoing operations, such as computing, communication, and disk loading. Since only a small segment (whose size is referred to as the layer window size hereafter) is reloaded per round, memory overflow can be avoided, and prefetched layers are less likely to be evicted, thus mitigating the prefetch-release conflict (see Figs. 4 and 5 in Appendix A.2 for examples). In addition, PRP processes both input and output on the head device, providing enhanced interaction privacy.
@@ 3.1 Pipelined-Ring Parallelism with Prefetching
Figure 1: Pipelined-ring parallelism. In this case, 6 devices handle a 36-layer model. With a layer window size of 2, the model is split into 18 segments and assigned to 6 devices in ring order, so each device needs 3 rounds to predict one token.
In Fig. 1, all devices share a layer window size of 2. Each device processes two model layers per round and takes 3 rounds to predict one token. Fig. 6 a,b in Appendix A.2 show the PRP timeline on homogeneous devices with fast and slow disks. On fast disks, prefetching latency is fully overlapped; on slow disks, CPU and disk operate alternately, removing bubbles but leaving residual page-fault loading latency, since required layers may not be fully loaded when computation begins. Fig. 2 evaluates this multi-round design: for large models, PRP reduces PP’s TPOT by about 50% (PP is equivalent to PRP at $k=1$), while for small models, PRP converges to PP with similar TPOT.
In practice, however, home devices are heterogeneous. Using a uniform layer window size still produces pipeline bubbles (see Fig. 6c in Appendix A.2), whereas assigning different window sizes per device can reduce them (see Fig. 6f in Appendix A.2). In general, stronger devices should be assigned larger window sizes. However, determining which devices are stronger and how large their window sizes should be is challenging. Beyond common heterogeneity in computing, communication, and memory, disk loading is heavily affected by OS-specific memory reclamation and disk read throughput. Such complex heterogeneity makes disk latency difficult to quantify.